OpenStudy (anonymous):

Solve 4x^2 = 12x + 40.

OpenStudy (anonymous):

I solved it, and i'm not sure if it's right or not.. I got x = -2 and x = -5

OpenStudy (vishweshshrimali5):

x^2 - 3x - 10 = 0 => (x-5)(x+2) = 0 => x = 5 and x = -2

OpenStudy (anonymous):

\[ 4x^{2} - 12x - 40 = 0 \] \[x^{2} - 3x - 10 = 0 \] Then Factor \[(x-5)(x+2) = 0\] So x = -2 and +5, you got it.

OpenStudy (anonymous):

Thank you! (:

OpenStudy (vishweshshrimali5):

ur welcome